At first an athlete jogs at 4 miles per hour and then jogs at 8 miles per hour, traveling 6 miles in 1 hour. How long does the athlete jog at each speed?
Answer by ankor@dixie-net.com(22740)

At first an athlete jogs at 4 miles per hour and then jogs at 8 miles per hour, traveling 6 miles in 1 hour.
How long does the athlete jog at each speed?
:
let t = time jogging at 8 mph
then
(1-t) = time jogging at 4 mph
:
Write a distance equation: dist = speed * time
8mph dist + 4mph dist = 6 mi
8t + 4(1-t) = 6
8t + 4 - 4t = 6
8t - 4t = 6 - 4
4t = 2
t = 2/4
t = 1/2 hr at 8 mph, then obviously, 1/2 hr at 4 mph also